OpenStudy (anonymous):

What is the names of these proof properties? Please help! 1. AB+BC=AC 2. 3n+3(n+4)=42 3. 3n+3n+12=42 4. 6n+12=42 5. 6n=30 6. n=5 word bank- simplify angle addition segment addition distributive addition addition property subtraction property division property multiplication property substitution

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@SasukeSai<3

OpenStudy (perl):

1. segment addition

OpenStudy (anonymous):

Thanks, #2?

OpenStudy (perl):

not sure about 2. i would need more context

OpenStudy (perl):

is there a proof ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

2. distributive property

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Do you know any others?

OpenStudy (anonymous):

i'll see

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

i think 3 is angle addition.. im not very sure but sure

OpenStudy (anonymous):

Okay, well that's more than what I have lol

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

either number 5 or 6 is substitution for sure

OpenStudy (anonymous):

im just not sure which one

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

but i dont know number 4 sorry!

OpenStudy (anonymous):

Okay, that's fine, you really helped. Thank you so much :)

OpenStudy (anonymous):

your welcome! no problem!

OpenStudy (anonymous):

@perl do you know what 4, 5, & 6 is?

OpenStudy (perl):

yes

OpenStudy (anonymous):

Okay thanks, could you help me?

OpenStudy (perl):

What is the names of these proof properties? Please help! 1. AB+BC=AC segment addition 2. 3n+3(n+4)=42 not sure 3. 3n+3n+12=42 distribute property 4. 6n+12=42 addition property 5. 6n=30 subtract property 6. n=5 division property word bank- simplify angle addition segment addition distributive addition addition property subtraction property division property multiplication property substitution

OpenStudy (anonymous):

Thank you so much <3